OpenStudy (anonymous):

can x<-2 U x>2 be solved?

OpenStudy (anonymous):

x2 – 3 = 2x x2 – 2x – 3 = 0 (x – 3)(x + 1) = 0 x – 3 = 0 or x + 1 = 0 x = 3 or x = –1 Then the solution to x2 – 3 = 2x is x = –1, 3

OpenStudy (anonymous):

okay, because the question asks to pick a graph, one wilth open circles on -2 and 2 pointing out, which i think is the answer,

hartnn (hartnn):

it can be plotted,as its the inequality,like this